avakis FN 305-Foundations of Finandal Management Soring 2019- Final Exam Student: Directions: Complete all answers on separate paper. Use pencil for all problems. Show the formula first and can check your answer with the financial calculator, but do not just write down a number you got from your calculator. The problems count 23 points each and the discusslon question counts 39 points for a total of 200 points. Use a pen to write the discussion question. (Assume 365 days per year for all your calculations. You calculations requiring daily interest rates.) Problems (23 points each) 1. Your father has promised to give you $500,000 when you graduate with your master's degree in international business from California International University. If you plan to graduate in exactly 5 years from now, how much money should he put in a savings account today to ensure he has the money to give you? Assume the interest rates will remain constant at 3.33 % 2. You like to gamble so you play the stock market! You bought 100 shares of Disney Company (they make movies) for a price of $90 per share 4 years ago. Now, the price is $140 per share and you decided to sell all your shares. During this time, you collected dividends of $1.75 per share each year. How much in dollar terms was your profit? What was your return on investment in per annum terms expressed as a percent? 3. You received your credit card statement and the company had added interest to your account from last month of $7.26 because you only paid part of the bill. The part you did not pay was equal to $740. If your billing cycle was 31 days last month and you did not make any new purchases, what annual interest rate are they charging you on your account? 4. You want to buy a new car. The price of the car is $35,000 but your rich uncle is going to pay the down payment for you which is 10 % of the sale price. You have to take a loan for the remainder. If the bank offers you a rate of 4.125% and you finance the loan for 5 years, what will be your monthly payments on the loan
Discussion (39 points) In a good paragraph with at least 10 sentences, discuss the concept of "capital budgeting." Explain at least three things about this concept: what is it used for in corporations, why does the method insure against future loss, where does the interest rate (WACC) used in calculations come from and how does it differ from the internal rate of return (IRR) on the project itself. Writing skills will be evaluated as well as your knowledge of the concepts and your ability to explain them. Assume you are explaining these ideas to someone who has never had any course in finance. Don't use any technical terms without explaining what they mean in finance. Formulas 1 - (1+) PVA = PMT FV= PV (1+i)" FV= PV ert [(1 +i) FVA = PMT FV PV= u(1+T) Selling Price - Original Cost RR (ROI) Original Cost CFn CF2 NPV= CFo +1+r)! * (1+r}2 CF + + + (1+r) CFt n u+T 0-17
4. You want to buy a new car. The price of the car is $35,000 but your rich uncle is going to pay the down payment for you which is 10 % of the sale price. You have to take a loan for the remainder. If the bank offers you a rate of 4.125% and you finance the loan for 5 years, what will be your monthly payments on the loan? 5. You plan to retire in 45 years with $3,500,000 in the bank. If the interest rate is constant at 4.2 %, how much will you have to save and put into your 401 K account every month to reach your goal? 6. Your bank calculates interest on savings accounts on a continuous basis. If you deposit $6,000 in a savings account and the interest rate is 2.75 % , what will be the total amount in your account after 4 years. Assue additional deposits were made. 7. You are working for a car park company and they are considering construction of a new wing to their commercial parking garage. The estimated cost of construction of the new wing to the old building is $40,000 which has to be paid at the beginning of the project. The cash receipts from customers who will park in the new wing are estimated to be as follows for the next 5 years: Cash Flow End of Year $ 3,000 $ 8,000 $13,000 $14,500 $ 22,000 If you have a Financial Calculator, please calculate the IRR as well and use this in your decision. 2 4 5 What is the Net Present Value to the company if they build this new garage? Assume the company has a weighted average cost of capital of 9 %. Do you recommend the company adopt this project? Why?
Answer #1

Question 1:

Here, we are trying to calculate the present value which is the amount of money he should put into the savings account today

FV = 500,000

r = 3.33% = 0.0333

n =5

PV = ?

PV = FV/(1+r)^n

PV = 500,000/(1+0.0333)^5

PV = 500,000/1.0333^5

PV = $424,461.06

So, your father must put in $424,461.06 into the savings account today

Note: We have answered one full question. Please note that only one full question will be answered at a time. Please post each question separately for experts to answer